if 9 th term of an A.P. is zero,prove that 29 th term is double than the 19 th term

if 9th term of an A.P. is zero,prove that 29th term is double than the 19th term

Grade:11

1 Answers

Mohd Shoaib
6 years ago
As a9 = o So, a + 8d = 0 a = -8d............ ...(i)Now , a29 = a + 28d a29 = -8d + 28da29 = 20d................(ii)a19 = a+ 18da19 = -8d + 18da19 = 10d.................(iii)Now from eq (ii) and (iii) 20d = 2* 10 da29 = 2*a19Hence proved......